OpenStudy (anonymous):

Solve each equation. Check your solution. 10. 3(g-3)=6 12. 5(2c+7)=80 14. 3(a-3)=2(a+4) 16. 4(x-2)=3(1.5+x)

OpenStudy (anonymous):

I'll show you how to do the first one then you can try the next one :o) 3(g-3)=6 3g-9=6 +9 +9 3g=15 /3 /3 g=5 Can you do the next one? :o)

OpenStudy (anonymous):

Yeah, I'm gonna write that one down real quick. If I don't get the second one can you help me out? :)

OpenStudy (anonymous):

Of course! :o)

OpenStudy (anonymous):

Ok, thank you! One minute

OpenStudy (anonymous):

5(2c+7)=80 10c+35=80

OpenStudy (anonymous):

I don't know if I subtract 35 or subtract 80

OpenStudy (anonymous):

Awesome job!!! And you're going to subtract 35 from both sides so that you can isolate c. :o)

OpenStudy (anonymous):

10c+35=80 -35 -35 10c=45 /10 c=4.5

OpenStudy (anonymous):

Perfect, spot on! Try the next one, you can do it. :o)

OpenStudy (anonymous):

Ok thank you! :D The assignment is 10-30 even and then a few more after that, can I tell you them when I'm done and have you check them?

OpenStudy (anonymous):

Of course :o) I know you're gonna do great!

OpenStudy (anonymous):

You're so nice! Thank you! I needed help really bad and nobody would help but I'm glad you are!

OpenStudy (anonymous):

14. 3(a-3)=2(a+4) 3a-9=2a+8

OpenStudy (anonymous):

Do I subtract from 3a or 2a

OpenStudy (anonymous):

If you ever need help you can post a question and tag me or you can message it to me, I'll always help the best I can :o) Perfect again! Let's see, I don't think it matters which way you do it. But it'd probably be easier if you did: 3a-9=2a+8 -2a -2a

OpenStudy (anonymous):

Ok then it would be 1a-9=8 -8 -8

OpenStudy (anonymous):

Then it would be 1a=-17 and then divide by 1?

OpenStudy (anonymous):

Remember 1a can be represented as just a :o) So a=-17 :o)

OpenStudy (anonymous):

However, if you did divide by 1 you'd still get a=-17 :o)

OpenStudy (anonymous):

Ohh, so it's a=-17?

OpenStudy (anonymous):

Nope. Lets go back a few steps. Can you see where you went wrong with this step? 1a-9=8 -8 -8

OpenStudy (anonymous):

Would it be just 9-8? Cause I think I took -9-8

OpenStudy (anonymous):

Well, if you subtract 8 we wouldn't have anything left on the other side. So instead of subtracting 8, we're actually going to add 9 to both sides like this: 1a-9=8 +9 +9

OpenStudy (anonymous):

a=17?

OpenStudy (anonymous):

Hang in there! Remember when you're doing this, you're always going to take from the side with the variable. You're trying to get the variable on the side of the equation alone. So, you're taking the numbers from that side to get a alone. Bingo! :o)

OpenStudy (anonymous):

Yay :D

OpenStudy (anonymous):

4(x-2)=3(1.5+x) 4x-8=4.5+3x +8 +8

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

You have 4x-8=4.5+3x correct, but do you notice how there is an x on both sides? Remember we're trying to get the x's on one side of the equal sign, not both. Keeping this in mind, what do you think you're first step will be?

OpenStudy (anonymous):

Oh yeahh! I knew that. It would be -4x from both sides?

OpenStudy (anonymous):

Well, you have +4x and +3x, do you think it'd be easier to do +4x-3x or +3x-4x? :o)

OpenStudy (anonymous):

-3x?

OpenStudy (anonymous):

Right! You could think of it like, would you rather do 4-3 or 3-4? So, 4x-8=4.5+3x -3x -3x

OpenStudy (anonymous):

then it would be 1x-8=4.5 +8 +8

OpenStudy (anonymous):

You got it! Awesome! So whats the answer?

OpenStudy (anonymous):

x=12.5

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Awesome work, so proud of you!! :o)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

If you have anymore questions or get stuck when doing them lemme know and I'll help :o)

OpenStudy (anonymous):

Ok :D Thank you! I'm gonna start on the next one :)

OpenStudy (anonymous):

2(3.5n+6)=2.5n-2 7n+12=2.5n-2 -2.5n -2.5n ---------------- 4.5n+12=-2

OpenStudy (anonymous):

Perfect, keep going :o)

OpenStudy (anonymous):

n=-3.1

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

Omg :D

OpenStudy (anonymous):

4(f+3)+5=17+4f 4f+12+5=17+4f -4f -4f ------------------ 0f+12+5=17

OpenStudy (anonymous):

This looks like a strange one! Well, it looks like you did it correct though. :o) So, you can see that f=0 already because of the 0 in front of f. If there are 0 f's, you really have no f's at all. So, there really is no point in solving it farther. :o)

OpenStudy (anonymous):

Oh, my teacher said something about "Many Solutions" and "No Solution"

OpenStudy (anonymous):

You can check it to by plugging 0 in for f: 4f+12+5=17+4f 4(0)+12+5=17+4(0) 0+12+5=17+0 17=17✓ It's probably "No Solution" then. :o)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You can check all your other answers to by plugging in your answer into the original equation like I just showed you. If they equal each other, it's correct. If they don't, then your answer is wrong. Anytime! :o) Got anymore? :o)

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

You're not, don't worry about it! If you're not sure or want me to check it, just post it. :o)

OpenStudy (anonymous):

8y-3=5(y-1)+3y 8y-3=5y-5+3y -5y -5y --------------- 3y-3=-5+3y

OpenStudy (anonymous):

Right? so far?

OpenStudy (anonymous):

Well, notice the 5y and 3y on the same side? You can add these like terms together and then put them on the other side. 8y-3=5y-5+3y 8y-3=8y-5 Looks like the same thing as the last problem.

OpenStudy (anonymous):

However, how you did it is correct as well, you'd end up with the same thing. :o)

OpenStudy (anonymous):

I get it! :D

OpenStudy (anonymous):

Great! :D

OpenStudy (anonymous):

-8y from both sides? then 0-3=-5

OpenStudy (anonymous):

Yes ma'am :o)

OpenStudy (anonymous):

Then what would I do next? Sorry my brain stopped working. xD

OpenStudy (anonymous):

Well, just like the other one with f, we can already see that y=0 and there is no solution. :o)

OpenStudy (anonymous):

Ohh ok!

OpenStudy (anonymous):

1/2(2n-5)=4n-1

OpenStudy (anonymous):

Okay, so 1/2 = .5. So 2*.5 ( the same as half) is 1. -5*.5= -2.5 n-2.5=4n-1

OpenStudy (anonymous):

then -4n on both sides ------------- -3-2.5=-1

OpenStudy (anonymous):

Nope, we will take n from the lest side: -2.5=3n-1

OpenStudy (anonymous):

Ohh!

OpenStudy (anonymous):

You hanging in there? :o) You're doing great!

OpenStudy (anonymous):

Yeah I'm trying :)

OpenStudy (anonymous):

then add 2.5?

OpenStudy (anonymous):

Close, we're going to add 1. Is it confusing you cause its on the opposite side? Is this better?: 3n-1=-2.5

OpenStudy (anonymous):

Yeah but i have it wrote down like -2.5=3n-1

OpenStudy (anonymous):

n=-0.5

OpenStudy (anonymous):

Hey, can you help me with 2 more and then the rest tomorrow?

OpenStudy (anonymous):

Sure :o) And yup, n=-0.5, great job! :o)

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

8y-3=5(y-1)+3y

OpenStudy (anonymous):

No solution

OpenStudy (anonymous):

Fantastic! :o)

OpenStudy (anonymous):

-3(4b-10)=1/2(-24b+60)

OpenStudy (anonymous):

Noo solution

OpenStudy (anonymous):

Muy bien! Correcto! :o)

OpenStudy (anonymous):

omg :D

OpenStudy (anonymous):

The next one I need help on

OpenStudy (anonymous):

Alright, let's see it :o)

OpenStudy (anonymous):

d/0.4=2d+1.24

OpenStudy (anonymous):

\[\frac{ d }{ 0.4 }=2d+1.24\]Okay, do you know how to cancel the d when its dividing? :o)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Alright, well when we were doing it before, you notice how we were subtracting if the number was adding? We're going to do the same thing here. What's the opposite of dividing?

OpenStudy (anonymous):

multiplying!

OpenStudy (anonymous):

Bingo! So, we're going to multiply d on both sides. Can you do it? :o)

OpenStudy (anonymous):

Wait, so I divide 0.4?

OpenStudy (anonymous):

?